An open-topped cylindrical pot is to have volume 55 cm3. Determine the minimum possible amount of material used in making this pot? Show your complete solution
Answer by Boreal(15235)

The area is πr^2+2πrh
so πr(r+2h)=A
V=πr^2h=55,
so h=55/πr^2=17.507/r^2
A=πr(r+(110/πr^2))=πr^2+(110/r)
take the derivative, and set that equal to 0.
2πr-110/r^2=0
multiply through by r^2 and move the second term to the right
2πr^3=110
πr^3=55
r^3=17.507
r=2.60 cm
h=17.507cm^3/(2.60 cm)^2
=2.60 cm
So the total surface area is πr(r+2h)=2.60π(7.80), with units cm*cm or cm^2=63.54 cm^2 (rounding at end)